Let f(x)=x^2+7 and g(x)=x^2+4. Find the following functions, and simplify each as much as possible: your final answer should be a polynomial with only one term in each power of x.
f(x)+g(x)=
f(x)-g(x)=
f(x)g(x)=
f(x)/g(x)=

Answer by Cromlix(4381)

Hi there,
Let f(x)=x^2+7 and g(x)=x^2+4.
f(x)+g(x)= 2x^2 + 11
f(x)-g(x)= 3
f(x)g(x)= x^4 + 11x^2 + 28
f(x)/g(x)= x^2+7/x^2+4
Hope this helps :-)